Hacker Software has 11.6 percent coupon bonds on the market with 14 years to maturity. The bonds make semiannual payments and currently sell for 108.6 percent of par.

What is the current yield on the bonds? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Current yield %  

What is the YTM? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  YTM %  

What is the effective annual yield? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

  Effective annual yield %  
Answer #1

Answer:

Current yield:

= Annual coupon payment÷Bond price

= $11.6÷$108.6

= 10.68%

Yield to maturity:

PV = pmt(1-(1+r)-n) / r + FV(1+r)-n
1086=58(1-(1+r)-28) / r + 1000 x (1+r)-28
Solving for r we get r =5.21%
Annual YM = 5.21% x 2 = 10.42%

Effective Annual Yield

=(1+r)^m-1
=(1+0.0521)^2 -1
=1.1069-1
=10.69%